OpenStudy (anonymous):

a damaged carpet which cpst $180 when new, is sold for $100 what is the percentage loss?

OpenStudy (anonymous):

loss%=loss/c.p. *100=80/100 *100 =80%

OpenStudy (anonymous):

sorry ,80/180 *100 =

OpenStudy (anonymous):

44.44%

OpenStudy (anonymous):

\[\LARGE 180(1-x\%)=100\] \[\LARGE 180-\frac{180x}{100}=100\] \[\LARGE -\frac{18x}{10}=100-180\] \[\LARGE -\frac{18x}{10}=-80 \] \[\LARGE \frac{18x}{10}=80 \] \[\LARGE 18x=800 \] \[\LARGE x=\frac{800}{18} \] \[\LARGE x\approx 44 \]

OpenStudy (anonymous):

whoa kresh !!!where do you get that style???